4. For a solution of Cys (C), at what pH would the net charge be closest to 1? a) 10.7 b) 8.4 c) (1.9-8.4)/2 = 5.5 d) (10.7 + 8.4)/2 = 9.55 5. For a solution of Tyr (Y), what percentage of the side chain R groups OH would be deprotonated at pH = 10.5? a) 0% b) 25% c) 50% d) 100% 6. For the peptide His -Asp-Gln-Glu-Ala- Lys (HDQEAK), what would be the closest overall charge at pH 8.0? the closest overall charge at pH8.0 a)0 b)+ c)+2 d)-1 c)-2 7. For the same peptide as in Question 6, His-Asp-Gln-Glu-Ala-Lys (= HDQEAK), what would be the closest overall charge at pH 1 ? a)0 b)+2 c)+3 d) c)-2

Can someone explain how these work?
Answer #1

4) pI of cysteine is 5.07. So, at pH greater than 5.07, charge will be negative. In these options, 5.5 (option c) is is the closest option. So, 'c' option is the correct answer.

5) pKa = pH + log [protonated form]/[deprotonated form]

10.5-10.5 = log [protonated form]/[deprotonated form]

10^0 = [protonated form]/[deprotonated form]=1

So, [protonated form] = [deprotonated form] (So, 50% will be protonated and the rest 50% will be deprotonated.)

option c. is correct.